OpenStudy (anonymous):

Solve by factoring x^2-6x-7=0

OpenStudy (anonymous):

x^2-7x+x-7

OpenStudy (anonymous):

(x-7)(x+1) are you having difficulty with factoring?

OpenStudy (anonymous):

Yes .

OpenStudy (anonymous):

So x(x-7)+1(x-7)=0 (x-7)(x+1)=0 x=7,-1

OpenStudy (anonymous):

what is it that youre having trouble with?

OpenStudy (anonymous):

- I just dont understand it . I dont know what it is I have to do .

OpenStudy (anonymous):

Allrite here we go then Can u see the last term and its sign

OpenStudy (anonymous):

you need to find two numbers that add up to your middle number and multiply to the last in this problem you need two numbers that add to -6 and multiply to -7

OpenStudy (anonymous):

first you fin the factors of seven, the only thing it could be is 1 and 7

OpenStudy (anonymous):

since seven is negative however, one of these values must also be negative (negative time positive equals negative)

OpenStudy (anonymous):

now look at the second number, -6, since its negative, the larger of the two factors (1 and 7) must be negative (-7+1=-6)

OpenStudy (anonymous):

so youre two factors are -7 and 1